Ben has $2.25 in nickels and dimes. He has 6 more nickels than dimes. How many nickels does he have?
I've tried everything I know, but I cant seem to solve it :(

Answer by scott8148(6628)

"Ben has $2.25 in nickels and dimes" ___ 5n + 10d = 225

"He has 6 more nickels than dimes" ___ n = d + 6

substituting ___ 5(d + 6) + 10d = 225 ___ 5d + 30 + 10d = 225 ___ 15d = 195

solve for d, then substitute back to find n